OpenStudy (anonymous):

i am working on soloving a system of equations though elimination can some one help me??

OpenStudy (anonymous):

3. What is a possible first step to eliminate a variable in the following system? 7x - 9y = 2 3x - 6y = -4 Multiply the first equation by 2 and the second equation by -3. Multiply the first equation by -3 and the second equation by -2. Multiply the first equation by 7 and the second equation by 3. Multiply the first equation by -2 and the second equation by 4.

OpenStudy (anonymous):

4. Solve 2x + y = 7 x - y = -4 (1, 5) (-1, 5) (1, -5) (-1, -5)

OpenStudy (anonymous):

5. Solve x + 2y = 13 4x + 8y = -9 (3, 5) (-1, 7) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

6. Solve 3x - 5y = 17 9x - 15y = -4 (-1, -4) (4, -1) No Solutions Infinitely Many Solutions

OpenStudy (anonymous):

7. Solve 5x + 2y = 7 10x + 4y = 14 (1, 1) (2, negative three halves) No Solution Infinitely Many Solutions

OpenStudy (anonymous):

8. Solve 7x - 2y = -3 14x + y = 14 (4, five sevenths) (4, seven fifths) (five sevenths, 4) (seven fifths, 4)

OpenStudy (studygurl14):

http://www.purplemath.com/modules/systlin5.htm

OpenStudy (anonymous):

LOL these many questions at once

OpenStudy (anonymous):

sorry but thats all

OpenStudy (studygurl14):

I could definitely help you with this. but it would take too long and I don't have a lot of time right now. sorry

OpenStudy (anonymous):

uggg :(

OpenStudy (anonymous):

First one if we see the options he is asking about the first step of the solution

OpenStudy (anonymous):

Lets verify with the options given and see which one can result us in elimination of a term

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I could sense its only the first option which satisfies elimination step so I would go with first one

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

42x + y = 7 x - y = -4 Adding these two equations we get ............................. 3x = 3 => x = 1 and now substitute x = 1 in any one of the equation to get y 1 - y = -4 => -y = -4 -1 => -y = -5 => y = 5

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

x + 2y = 13 4x + 8y = -9 Multiply first equation with 4 u get 4x + 8y = 52 4x + 8y = -9 Here both x and y will be eliminated if we subtract these two equations so there are no solutions for this

OpenStudy (anonymous):

Hope U can try the other one in same manner

OpenStudy (anonymous):

huh?? im lost

OpenStudy (anonymous):

For fourth one it is (1,5) and for fifth one Infinite solutions

OpenStudy (anonymous):

okay do u kno 678?

OpenStudy (anonymous):

6 , 7 also no solutions

OpenStudy (anonymous):

8th one u can solve for solutions

OpenStudy (anonymous):

Solve 7x - 2y = -3 14x + y = 14 Multiply equation 1 with 2 14x - 4y = -6 14x + y = 14 Subtracting both the equations we get 14x - 4y = -6 -14x -y = - 14 ................................... -5y = -20 => y = 4 Substitute y value in any of the equation 7x - 8 = -3 => 7x = 5 =>x = 5/7 (5/7 , 4) is ur answer

OpenStudy (anonymous):

thanks so so much!